# Improving Object Detection by Modifying Synthetic Data with Explainable AI

### v1

@@ Abstract
In many computer vision domains the collection of sufficient real-world data is challenging and can severely impact model performance, particularly when running inference on samples that are unseen or underrepresented in training. Synthetically generated images provide a promising solution, but it remains unclear how to design synthetic data to optimally improve model performance, for example whether to introduce more realism or more abstraction in such datasets. Here we propose a novel conceptual approach to improve the performance of computer vision models trained on synthetic images, by using robust Explainable AI (XAI) techniques to guide the modification of 3D models used to generate these images. Importantly, this framework allows both modifications that increase and decrease realism in synthetic data, which can both improve model performance. We illustrate this concept using a real-world example where data are sparse; the detection of vehicles in infrared imagery. We fine-tune an initial YOLOv8 model on the ATR DSIAC infrared dataset and synthetic images generated from 3D mesh models in the Unity gaming engine, and then use XAI saliency maps to guide modification of our Unity models. We show that synthetic data can improve detection of vehicles in orientations unseen in training by 4.6% (to mAP50 scores of 94.6%). We further improve performance by an additional 1.5% (to 96.1%) through our new XAI-guided approach, which reduces misclassifications through both increasing and decreasing the realism of different parts of the synthetic data. These proof-of-concept results pave the way for fine, XAI-controlled curation of synthetic datasets through detailed feature modifications, tailored to improve object detection performance.
@@ 1 Introduction
In many machine learning domains the collection of sufficient real-world data is challenging and can severely impact model performance, particularly when running inference on test samples and scenarios that are unseen or strongly under-represented in training data. Barriers to data collection often include time, cost, safety, complexity, and legal or privacy restrictions [1], and the defence [2], healthcare [3], finance [4], and autonomous systems [5] domains frequently face strict regulations that limit the sharing of sensitive data. For computer vision models, synthetically generated images provide a promising solution to overcome these constraints [6], and synthetic data can be generated by a range of techniques including synthetic composite imagery[7, 8, 9], GANs [10, 11, 12, 13], Diffusion models [14, 15], GAN-Diffusion hybrids [16] and 3D rendered environments [17, 18, 19].
@@
However, despite significant progress in the generation of synthetic data, it remains unclear how best to design synthetic data to optimally improve model performance, i.e. what constitutes the most useful or optimal synthetic data to augment a specific real-world dataset for any given task. For example, while synthetic data are associated with a number of issues such as the domain gap between real and synthetic samples [20, 21], poor generalisability owing to a lack of diversity [22, 23, 24], and mode collapse [25, 26, 27], approaches to mitigate against these issues range from building models aimed at producing more photorealistic outputs [28, 29], to methods that make data less realistic in order to disrupt sources of misclassification and force a network to learn the essential features of an object, such as domain randomisation [23]. Therefore even at a fundamental level, it is unclear when more realism or when less realism would be more beneficial in synthetic data.
One approach to address this issue is to let model performance feed back into the synthetic data generation process and directly guide modifications to the synthetic data, which could include both increases or decreases in data realism. In this paper, we propose a novel conceptual approach to achieve this by combining gaming engines, which are already extensively used for generating synthetic data [30, 31, 32, 33, 34, 35], with Explainable AI (XAI) techniques. Synthetic data have previously been used to improve computer vision performance [33, 34, 35]. There has also been effort on improving the robustness of computer vision explainability by grounding it in axiomatic approaches [36, 37]. We go one step further by combining these two methods. We use robust saliency maps to guide modification of 3D mesh models within a gaming engine (Unity3D [38]), in order to improve performance of an object detection model trained with synthetic data. Saliency-guided data augmentation was first used to mix and modify training samples [39], in order to preserve salient features of the samples being mixed, and therefore improve regularization by guiding the model to focus on less discriminative features. In our work, we adopt a fundamentally different approach by directly modifying mesh models used to generate synthetic data, guided by a more robust explainability technique (SHAP [40]), combined with a procedure to track and interpret how such modifications affect the model focus, i.e. which features in the input image contribute most to the model decision. A close analogue to our proposed method is the feature steering approach which identifies specific features using mechanistic interpretability [41], and weakens or strengthens their activation to influence model outputs. However this approach [42] does not change the underlying model weights and requires access to the complete model, whilst our method is model-agnostic and instead improves performance by steering model weights using targeted synthetic data modifications.
@@
In this paper we demonstrate our new approach using a real-world example where collection of data is challenging and where synthetic data are of great benefit for detection of unseen scenarios; the detection of vehicles in infrared data. We use the Defense Systems Information Analysis Centre Automated Target Recognition (DSIAC ATR) dataset and test how our XAI-guided approach to synthetic data modification can improve detection of vehicles in orientations that are unseen in real-world training data.
Our contributions are:
@@
  • We propose a novel method to use explainable-AI techniques to guide modification of 3D mesh models within a gaming engine, in order to improve performance of an object detection model trained with synthetic data.
@@
  • We show that XAI-guided modification of synthetic data, which involves both increasing and decreasing data realism, improves performance by 1.5% over an initial 4.6% gain from using a base version of synthetic data.
In this paper we first explain our conceptual approach in Section 2, and then introduce the DSIAC ATR dataset and explain how we generate synthetic data with the Unity gaming engine in Section 3. In Section 4 we explain our experimental setup. Finally, we present our results in Section 5 and discuss the broader implications of this work in Section 6.
@@ 2 Method
We propose a method to improve the performance of object detection and classification algorithms that are trained on synthetic images, through using saliency maps to guide modification of 3D mesh models used to generate the synthetic data. We first motivate our approach with a mathematical toy model that illustrates how characteristics of training data can affect model solutions. In particular, we introduce and explain the common features and unique features for a pair of classes, which are key to understanding the theoretical rationale for our methodology.
@@
For our toy model, we consider an MLP binary classifier for classes $k$ and $k^{\prime}$, that decomposes an input sample $\mathbf{x}_{j}$ from class $j\in\{k,k^{\prime}\}$ as a composition of $N$ features, $\mathbf{v}_{1},\ldots,\mathbf{v}_{N}$, encoded by $N$ neurons of the hidden layer assuming monosemanticity [43], and assigns a confidence score for class $k$ via the final classification layer as $f_{k}(\mathbf{x}_{j})\triangleq\sum_{i=1}^{N}w^{k}_{i}\phi(\mathbf{x}_{j},\mathbf{v}_{i})\triangleq\boldsymbol{\phi}(\mathbf{x}_{j})^{T}\mathbf{w}_{k}$. Here, $\phi(\mathbf{x}_{j},\mathbf{v}_{i})$ is a quantification of the extent of the presence of feature $\mathbf{v}_{i}$ in $\mathbf{x}_{j}$, encoded by the neuron activations in the hidden layer. This modelling of feature representations follows the framework introduced by OpenAI in [44], where specific features and the extents to which they are present in the input are represented by individual neurons and their activations, respectively, and provides justification for more general applicability of the assumption of monosemanticity, at least for higher level features. By training the classification layer $\mathbf{w}_{k}$ to simultaneously maximise the expected confidence of all correct classifications of class $k$, whilst also minimising the expected confidence for all misclassifications of $\mathbf{x}_{k^{\prime}}$ to class $k$, we get optimal weights across the whole dataset given by: $\mathbf{w}_{k}^{*}\propto\mathbb{E}[\boldsymbol{\phi}(\mathbf{x}_{k})]-\mathbb{E}[\boldsymbol{\phi}(\mathbf{x}_{k^{\prime}})].$
Therefore, in our trained binary classifier, large weights (large values in $\mathbf{w}_{k}^{*}$) are assigned to unique features for class $k$, those which are prominent in class $k$, i.e., features $\mathbf{v}_{i}$ for which $\mathbb{E}[\phi(\mathbf{x}_{k},\mathbf{v}_{i})]$ is high, but are less prominent in the other class $k^{\prime}$. Conversely, small weights (small values in $\mathbf{w}_{k}^{*}$) are assigned to common features, those that are prominent in both classes.
@@
As such, misclassifications between classes are more likely to arise when unique features have weights that are too small, and when common features have weights that are too large. Although this toy problem is simple, we suggest that this concept is more generally applicable to deep learning classification models. Our proposed procedure focuses on identifying unique and common features and then reinforcing and disrupting them respectively. In order to reduce confusion between two classes, we make modifications to the mesh models of these classes to promote unique features in each model (which we term Reinforcing modification) or to suppress common features associated with misclassification (which we term Disruptive modification).
@@
It is very important to note that although in this study we predominantly focus on reinforcing through increased realism and disruption through decreased realism, reinforcing modification could equally involve decreased realism (e.g. exaggeration of unique features) and disruptive modification could involve increased realism (e.g. greater, more realistic variety of common features).
@@ 2.1 Procedure
Our proposed procedure, which exploits this ability to reinforce unique features and disrupt common features to reduce misclassifications and improve model performance, is laid out in the following six steps. These same six steps are also illustrated in Fig. 1, for a simple graphical example of an object detection model trained on real and synthetic thermal infrared (MWIR) images of a hatchback car, a sports car and a truck.
  1. 1.
    Train an object detection model on the initial real and synthetic images.
  2. 2.
    Evaluate the performance of the model. We identify if the model performance is sufficient for our requirements. If it is we terminate the algorithm here.
  3. 3.
    We use the model confusion matrix to identify a target misclassification, typically the largest off-diagonal, non-background element (red square, Fig. 1), and the associated classes (blue=hatchback, green=sports car).
  4. 4.
    We identify the cause of the misclassification. We calculate SHAP saliency contributions (explained in more detail in the next section) for the correct classifications of both classes and for the misclassification. By examining correlations in the location, pattern and size of the SHAP values between the correct classifications and the misclassification, we identify unique features that help distinguish between the two classes and common features of the two classes that can lead to confusion.
  5. 5.
    We make modifications to the mesh models to reinforce unique features such that they become more prominent or disrupt common features such that they become less prominent.
  6. 6.
    The modified mesh models are used to generate a new version of the synthetic dataset, and we revert to step 1., where we retrain the model. This whole process is then repeated until sufficient model performance criteria have been met.
@@
We now provide additional details on our use of XAI in this procedure. We use SHAP [40] as an XAI tool because it is model-agnostic and satisfies desirable theoretical properties for explainability [45] over alternatives such as attention [46]. SHAP values [40], which are an approximation to Shapley values [47], take an input set of superpixels within the bounding box for each image and attribute the marginal contributions of each superpixel towards a specific classification. To calculate the superpixel attributions for each sample, a random mask is applied over the image, where the masked portions of the sample are replaced with samples from a background set. Inference is then computed over this masked image and each pixel is assigned its classification score. This is repeated for the same image over $1000$ different random masks to find the average contribution of a given pixel towards a specific classification, using the computationally efficient SHAP Kernel method [40].
We compute the pixel attributions over $50$ randomly selected test samples for which the model predicts high confidence for the specified class. For each pixel, we count the number of samples in which it has a contribution greater than a user-defined threshold (we use $40\%$ of the highest contribution in each sample, which represents a trade-off between information and noise). Thus we obtain an average SHAP contribution map within the bounding box, where each pixel is assigned a score of the number of samples in which it contributed more than the given threshold. In order to better visualise the average SHAP contribution maps, we further mask out the pixels that have a score less than another user-defined threshold (we use $50\%$ of the highest score), and superimpose it on a representative test image to identify the parts of the target contributing to the specific classification (see purple masks on Model Focus images in Fig. 5).
@@
Next we detail how we use SHAP contribution maps in practice to identify what area of the mesh model to modify, and to determine what type of change is needed (steps 4 and 5 in the previous section). We plot the average SHAP contribution maps inside the bounding boxes for samples in the two target classes (for the case class A is being misidentified as class B). Three different saliency maps are plotted; the saliency map for correct classification of class A, the saliency map for correct classification of class B, and the saliency map corresponding to the misclassification of class A as class B. Both unique and common features can be identified by comparison of these three SHAP saliency maps, with reference to the ground truth models. If the two SHAP maps for the misclassification (class A as class B) and correct classification of class B both have high saliency in the same location and the two datasets used to calculate these SHAP images are similar (i.e. have similar brightness, patterns) in this same region, then this is a common feature. Conversely, if a high saliency pattern at a location in the correct classification bounding box does not appear with any significant saliency in the misclassification at the same location, then this is a unique feature. It is important to note that this comparison must also account for variation across each class. For this reason we first need to cluster validation samples, using dimensional reduction techniques (such as principal component analysis), or using appropriate data labels corresponding to a natural dimension of variation (e.g. vehicle orientation in our example). The comparison of SHAP images is then carried out across each cluster.
@@ 3 Dataset
ATR DSIAC dataset. The ATR DSIAC dataset [52] was collected by the former US Army Night Vision and Electronic Sensors Directorate (NVESD). The dataset contains mid-wave infrared (MWIR, shortened to IR from here on) and electro-optical videos of $8$ different vehicle classes (Pickup Truck, SUV, BTR70, BRDM2, BMP2, T72, ZSU23-4, and 2S3) taken at daytime and night time at distances of $1000$m, $1500$m, and $2000$m (see Fig. 2a and supplementary Fig. 6 for examples of complete frames and closeups of all classes). In the videos the vehicles move in approximately circular paths, and the camera is located at a slightly elevated side-on position, so that a range of orientations (front, back, sides) are captured in the dataset. We extract video frames as image samples and only use night-time images to avoid rendering complexities caused by solar illumination and material reflectance when generating our synthetic dataset. For more details see supplementary material, Section 7.
Synthetic data generation. To generate synthetic data, we use the Unity3D gaming engine along with 3D models of the vehicle classes that are freely available online. We use these to develop a 3D scene with similar distance ranges and camera resolutions to the real IR images in the ATR DSIAC dataset. To simulate infrared image capture, we use the Probuilder tool [51] in Unity to set material properties (reflectance and emission) of the face of each mesh element, as illustrated by the selected orange panel in Fig. 3b.
We set the material properties and the scene details to produce renderings with roughly similar pixel values of the vehicles as in the original images from the dataset, and we implement a 2D convolution of the rendered image with an Airy-disk kernel to simulate the diffraction effect from a circular pixel aperture. In future it would be possible to automate this procedure, including setting of material properties and other parameters, but for this proof of concept we do this manually; a key benefit of our method is that it does not require photorealistic synthetic data due to the XAI-guided procedure, which focuses on performance gain from synthetic data alone, and therefore may modify synthetic data to be less realistic as well as more realistic.

### v2

@@ Abstract
Limited real-world data severely impacts model performance in many computer vision domains, particularly for samples that are underrepresented in training. Synthetically generated images are a promising solution, but 1) it remains unclear how to design synthetic training data to optimally improve model performance (e.g, whether and where to introduce more realism or more abstraction) and 2) the domain expertise, time and effort required from human operators for this design and optimisation process represents a major practical challenge. Here we propose a novel conceptual approach to improve the efficiency of designing synthetic images, by using robust Explainable AI (XAI) techniques to guide a human-in-the-loop process of modifying 3D mesh models used to generate these images. Importantly, this framework allows both modifications that increase and decrease realism in synthetic data, which can both improve model performance. We illustrate this concept using a real-world example where data are sparse; detection of vehicles in infrared imagery. We fine-tune an initial YOLOv8 model on the ATR DSIAC infrared dataset and synthetic images generated from 3D mesh models in the Unity gaming engine, and then use XAI saliency maps to guide modification of our Unity models. We show that synthetic data can improve detection of vehicles in orientations unseen in training by 4.6% (to mAP50 = 94.6%). We further improve performance by an additional 1.5% (to 96.1%) through our new XAI-guided approach, which reduces misclassifications through both increasing and decreasing the realism of different parts of the synthetic data. Our proof-of-concept results pave the way for fine, XAI-controlled curation of synthetic datasets tailored to improve object detection performance, whilst simultaneously reducing the burden on human operators in designing and optimising these datasets.
@@ 1 Introduction
In many machine learning domains the collection of sufficient real-world data is challenging and can severely impact model performance, particularly when running inference on test samples and scenarios that are unseen or strongly under-represented in training data. Barriers to data collection include time, cost, safety, complexity, and legal or privacy restrictions [55], and the defence [17], healthcare [13], finance [1], and autonomous systems [9] domains frequently face strict regulations that limit sharing of sensitive data. For computer vision models, synthetic training images provide a promising solution [8], and such data can be generated by a range of techniques including synthetic composite imagery [48, 60, 52], GANs [68, 7, 18, 20], Diffusion models [42, 46], GAN-Diffusion hybrids [63] and 3D rendered environments [40, 41, 5].
@@
However, despite significant progress in the generation of synthetic data, it remains unclear how best to design synthetic data to optimally improve model performance, i.e. what constitutes useful or optimal synthetic data to augment a real-world dataset for a given task. For example, synthetic data can be subject to a range of issues such as the domain gap between real and synthetic samples [3, 2], poor generalisability owing to lack of diversity [37, 57, 14], and mode collapse [54, 23, 21], but approaches to mitigate these issues range from increasing realism in synthetic data [34, 22], to making data less realistic in order to disrupt sources of misclassification and force a network to learn the essential features of an object (as in domain randomisation [57]). Therefore even at a fundamental level, it is unclear when more realism or when less realism is beneficial in synthetic data. Furthermore, whilst human-in-the-loop design and optimisation of synthetic datasets is desirable as it embeds human knowledge into the training process, improving model performance while also enhancing its robustness and generalisation [65], it requires significant domain expertise, time and effort from human operators. As these operators are already burdened with intense evaluation tasks (identifying failure modes, resolving them by annotating instances, curating the dataset, changing the training algorithm) this additional workload represents a major practical barrier to the optimal use of synthetic data [24].
One approach to simultaneously address both these issues is to let model performance directly tell the human operator where and what type of modifications are needed in the synthetic data, which could include increases or decreases in data realism as required. To this aim, we propose a novel explainable synthetic data generation framework with human-in-the-loop for object detection, using insights from Explainable AI (XAI) techniques to efficiently guide this process. Robust XAI techniques (SHAP [28]) are used to identify features that cause confusion or distinguish between classes, and the human operator then modifies the synthetic training dataset to respectively disrupt or reinforce these features, so as to steer the model’s attention and improve performance.
Related Work. Previous studies have developed techniques to test and ‘debug’ models by identifying low performing subsets of data using human-in-the-loop interactions [12, 11, 44, 45, 64], and then by fine-tuning models to improve their performance on these “hard samples”. In our approach, we instead use SHAP XAI [28] to identify features of the hard samples causing low model performance, and then carefully modify these features in the synthetic training data in order to improve performance on these hard samples.
Whilst previous works have have either improved computer vision performance using large volumes of synthetic data [47, 31, 51], or have investigated computer vision explainability [16, 25], our use of XAI to guide improvements in generation of new synthetic data is novel. Saliency-guided data augmentation has been used to mix and modify training samples [43], but we take a fundamentally different approach by directly modifying mesh models within a gaming engine (Unity3D [58]) that provides fine-grained control of textures, lighting and materials that, to the best of our knowledge, is not achievable with either this saliency-guided data augmentation [43] or with current controlled diffusion models [15, 27, 36].
A close analogue to our approach is “feature steering” which identifies specific features using mechanistic interpretability [4], and weakens or strengthens their activation to influence model outputs. However this approach [56] does not change the underlying model weights and requires access to the complete model, whilst our method is model-agnostic and instead improves performance by steering model weights using targeted modifications to synthetic data. Our method also differs to other methods such as using a varifocal loss [66] or online hard-example mining [53], by directly addressing the limitations of the dataset itself rather than optimising the training procedure alone.
@@
In this paper we use a real-world example of detection of vehicles in infrared data to demonstrate our new approach; for this task collection of data is challenging and synthetic data are of especial benefit for detection of unseen scenarios. We use the Defense Systems Information Analysis Centre Automated Target Recognition (DSIAC ATR) dataset to show how our XAI-guided approach to synthetic data modification can improve detection of vehicles in orientations unseen in real-world training data. Our contributions are:
@@
  • We propose a novel human-in-the-loop framework using explainable-AI techniques to guide modification of synthetic data, in order to improve performance of an object detection model trained with these data and reduce workload for the human operator.
@@
  • We show that XAI-guided modification of synthetic data, which involves both increasing and decreasing data realism, improves performance by up to 1.5% over an initial 4.6% gain from using a base version of synthetic data.
@@ 2 Method
We propose a method to improve the performance of object detection and classification algorithms that are trained on synthetic images, through using saliency maps to guide modification of these synthetic data. We first motivate our approach with a mathematical toy model that illustrates how characteristics of training data can affect model solutions. In particular, we introduce and explain the common features and unique features for a pair of classes, which are key to understanding the theoretical rationale for our methodology.
@@
For our toy model, we consider an MLP binary classifier for classes $k$ and $k^{\prime}$, that classifies an input sample $\mathbf{x}_{j}$, consisting of $N$ features, $\mathbf{v}_{1},\ldots,\mathbf{v}_{N}$, from class $j\in\{k,k^{\prime}\}$, and assigns a confidence score for class $k$ via a classification layer as $f_{k}(\mathbf{x}_{j})\triangleq\sum_{i=1}^{N}w^{k}_{i}\phi(\mathbf{x}_{j},\mathbf{v}_{i})\triangleq\boldsymbol{\phi}(\mathbf{x}_{j})^{T}\mathbf{w}_{k}$. Here, $\phi(\mathbf{x}_{j},\mathbf{v}_{i})$ is a quantification of the extent of the presence of feature $\mathbf{v}_{i}$ in $\mathbf{x}_{j}$. By training the classification layer $\mathbf{w}_{k}$ to simultaneously maximise the expected confidence of all correct classifications of class $k$, whilst also minimising the expected confidence for all misclassifications of $\mathbf{x}_{k^{\prime}}$ to class $k$, we get optimal weights across the whole dataset given by: $\mathbf{w}_{k}^{*}\propto\mathbb{E}[\boldsymbol{\phi}(\mathbf{x}_{k})]-\mathbb{E}[\boldsymbol{\phi}(\mathbf{x}_{k^{\prime}})].$
@@
Therefore, in our trained binary classifier, large weights (large values in $\mathbf{w}_{k}^{*}$) are assigned to unique features which are prominent in class $k$; i.e., features $\mathbf{v}_{i}$ for which $\mathbb{E}[\phi(\mathbf{x}_{k},\mathbf{v}_{i})]$ is high, but are less prominent in the other class $k^{\prime}$. Conversely, small weights (small values in $\mathbf{w}_{k}^{*}$) are assigned to common features, those that are prominent in both classes. As such, misclassifications between classes are more likely to arise when unique features have weights that are too small, and when common features have weights that are too large. Although this toy problem is simple, we suggest that this concept is more generally applicable to deep learning classification models. Our proposed procedure focuses on identifying unique and common features and then reinforcing and disrupting them respectively. In order to reduce confusion between two classes, we make modifications to the mesh models of these classes to promote unique features in each model (which we term Reinforcing modification) or to suppress common features associated with misclassification (which we term Disruptive modification).
@@
It is important to note that although here we predominantly focus on reinforcing through increased realism and disruption through decreased realism, reinforcing modification could equally involve decreased realism (e.g. exaggeration of unique features) and disruptive modification could involve increased realism (e.g. greater, more realistic variety of common features).
@@ 2.1 Procedure
Our proposed human-in-the-loop procedure, which exploits this ability to reinforce unique features and disrupt common features to reduce misclassifications and improve model performance, is laid out in the following six steps, which are also illustrated in Fig. 1 for a simple graphical example.
(1) Human operator trains an object detection model on initial real and synthetic images.
(2) Operator evaluates the performance of the model, and terminates procedure if it is sufficient for requirements.
(3) Operator identifies a target misclassification from the model confusion matrix, typically the largest off-diagonal, non-background element (red square in Fig. 1), and associated classes (blue=hatchback, green=sports car).
(4) Operator identifies cause of the misclassification, by calculating SHAP saliency contributions (more detail below) for the correct classifications of both classes and for the misclassification. Operator identifies unique features that distinguish between the two classes and common features that cause confusion, by examining correlations in the location, pattern and size of the SHAP values between the correct classifications and the misclassification.
(5) Operator makes modifications to the mesh models, guided by SHAP comparisons, to reinforce unique features or disrupt common features to respectively increase or decrease their prominence.
(6) Modified mesh models are used to generate new version of the synthetic dataset, and process reverts to step 1., where operator retrains the model. The process is repeated until sufficient model performance criteria are met.
@@
We now provide additional details on our use of XAI. We use SHAP [28] because it is model-agnostic and satisfies desirable theoretical properties for explainability [33] over alternatives such as attention [19]. SHAP values [28], which are an approximation to Shapley values [50], take an input set of superpixels within the bounding box for each image and attribute the marginal contributions of each superpixel to a specific classification. To calculate the superpixel attributions for each sample, a random mask is applied over the image, where masked portions of the sample are replaced with samples from a background set. Inference is then computed over this masked image and each pixel assigned its classification score. This is repeated for the same image over $1000$ different random masks to find the average contribution of a given pixel towards a specific classification, using the SHAP Kernel method [28]. We compute pixel attributions over $50$ randomly selected test samples for which the model predicts high confidence for the specified class. For each pixel, we count the number of samples in which it has a contribution greater than a user-defined threshold (we use $40\%$ of the highest contribution in each sample, balancing a trade-off between information and noise). Thus we obtain an average SHAP contribution map within the bounding box, where each pixel is assigned a score of the number of samples in which it contributed more than the given threshold. In order to better visualise the average SHAP contribution maps, we mask out pixels that have a score less than another user-defined threshold (we use $50\%$ of the highest score), and superimpose this on a representative test image to identify parts of the target that contribute most to the specific classification (see purple masks on Model Focus images in Fig. 5).
@@
Next we detail how we use SHAP contribution maps in practice to identify what area of the mesh model to modify, and to determine what type of change is needed (steps 4 and 5 in the previous section). We plot the average SHAP contribution maps inside the bounding boxes for samples in the two target classes (for the case where class A is misclassifed as class B). Three saliency maps are plotted; for the correct classification of class A, for the correct classification of class B, and the saliency map corresponding to the misclassification of class A as class B. Both unique and common features can be identified by comparison of these three SHAP saliency maps, with reference to the ground truth models. If the SHAP maps for the misclassification (class A as class B) and correct classification of class B both have high saliency in the same location and the two datasets used to calculate these SHAP images are similar (i.e. have similar brightness, patterns) in this same region, then this is a common feature. Conversely, if a high saliency pattern at a location in the correct classification bounding box does not appear with any significant saliency in the misclassification at the same location, then this is a unique feature. It is important to note that this comparison must also account for variation across each class. For this reason we first need to cluster validation samples, using dimensional reduction techniques (e.g. principal component analysis), or using appropriate data labels corresponding to a natural dimension of variation (e.g. vehicle orientation in our example). Comparison of SHAP images is then carried out across each cluster.
@@ 3 Dataset
The ATR DSIAC dataset [10] was collected by the former US Army Night Vision and Electronic Sensors Directorate (NVESD). The dataset contains mid-wave infrared (MWIR, shortened to IR from here on) and electro-optical videos of $8$ different vehicle classes (Pickup Truck, SUV, BTR70, BRDM2, BMP2, T72, ZSU23-4, and 2S3) taken at daytime and night time at distances of $1000$m, $1500$m, and $2000$m (see Fig. 2a and Fig. 1, Section 1 in supplementary material for examples of complete frames, closeups of all classes, and more details). In the videos the vehicles move in approximately circular paths, and the camera is located at a slightly elevated side-on position, so that a range of orientations (front, back, sides) are captured in the dataset. We extract video frames as image samples and only use night-time images to avoid rendering complexities caused by solar illumination and material reflectance when generating our synthetic dataset.
Synthetic data generation. To generate synthetic data, we use the Unity3D gaming engine along with 3D models of the vehicle classes that are freely available online. We use these to develop a 3D scene with similar distance ranges and camera resolutions to the real IR images in the ATR DSIAC dataset. To simulate infrared image capture, we use the Probuilder tool [39] in Unity to set material properties (reflectance and emission) of the face of each mesh element, as illustrated by the selected orange panel in Fig. 3b. We set the material properties and the scene details to produce renderings with roughly similar pixel values of the vehicles as in the original images from the dataset, and we implement a 2D convolution of the rendered image with an Airy-disk kernel to simulate the diffraction effect from a circular pixel aperture. In future it would be possible to automate this procedure, including setting of material properties and other parameters, but for this proof of concept we do this manually; a key benefit of our method is that it does not require photorealistic synthetic data due to the XAI-guided procedure, which focuses on performance gain from synthetic data alone, and therefore may modify synthetic data to be less realistic as well as more realistic.